john is two years less than four times as old as mary. john is also one year more than three times as old as mary. how old is each
Answer by stanbon(75887)

john is two years less than four times as old as mary. john is also one year more than three times as old as mary. how old is each
----
Equations:
j = 4m-2
j = 3m+1
----
Substitute and solve for "m":
4m-2 = 3m+1
m = 3 yrs (Mary's age now)
-------
Solve for "j":
j = 4m-2
j = 4*3-2 = 10 yrs (John's age now)
===================================
Cheers,
Stan H.